OpenStudy (anonymous):

Simple Question, but I don't get it at the moment -.- (4m-3)/(6m+8) = 3 => Solve the equation. (Pls. answer with explanation)

OpenStudy (anonymous):

(4m-3)/(6m+8) = 3/1 => (4m-3) = 3(6m+8) ....{ cross multiply } => 4m - 3 = 18m + 24 .. Can you solve it further .. ??

OpenStudy (anonymous):

Well, thank you! Solution is -27/14. Head was to full :D

OpenStudy (anonymous):

* -21/14 :P

OpenStudy (anonymous):

Good job! :)

OpenStudy (anonymous):

And, yeah the answer was -27/14 ..